OpenStudy (anonymous):

Anyone want to help with the integral of x/(x+2)^1/2 from -1 to 2?

OpenStudy (anonymous):

I just don't know how to set it up, what should u be?

OpenStudy (anonymous):

say u = 1/(sqrt(x-2) then you have integral of u/u^2-2. use partial fractions to solve.

OpenStudy (anonymous):

sorry,u = 1/sqrt(x+2)

OpenStudy (anonymous):

lemme know how it works out.

OpenStudy (anonymous):

ok thx I'll try it out

OpenStudy (anonymous):

did you get du=-1/2(x+2)^3/2? or did I do something wrong

OpenStudy (anonymous):

thats right, just checked it

OpenStudy (anonymous):

ok great! . i'd actually not done it. I just looked at the problem and did some calculation in my head. sorry.

OpenStudy (anonymous):

np, wait though not done yet

OpenStudy (anonymous):

so next du=-1(x+2)/2(x+2)^1/2 right?

OpenStudy (anonymous):

umm, you can solve it entirely in u and plug in the values of u, since this is a definite integral.

OpenStudy (anonymous):

no need to deal with x at all

OpenStudy (anonymous):

wait how come? I thought we had to make du look like the remaining terms in the integral, sorry I've just learned integrals today

OpenStudy (anonymous):

Okay. let me get back to you with this. I will take out my pen and paper and write it down. Do ask someone else too.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok, found it

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

Ok, so first Let \(u = (x+2)^{1/2} \implies du = (1/2)(x+2)^{-1/2}dx\) \[ \implies dx = 2(x+2)^{1/2} du\]

OpenStudy (anonymous):

So then \[\int_{-1}^2\frac{x}{(x+2)^{1/2}}dx = \int_1^2[u^2-2]du\]

OpenStudy (anonymous):

Whoops forgot the 2

OpenStudy (anonymous):

wow, ok let me try that out thanks, do u know of any rules to find these? or to you just have to guess the u sometimes?

OpenStudy (anonymous):

and does the integral change from -1 to 2, to 1 to 2, or is that a typo?

OpenStudy (anonymous):

The integral changes because \(u = (x+2)^{1/2}\) so when x = -1, u = 1 and when x = 2, u = 2

OpenStudy (anonymous):

right right, forgot about that, when you said you forgot the 2, is it ]2u^2-2]

OpenStudy (anonymous):

\(2[u^2 - 2]\)

OpenStudy (anonymous):

ah right, that would make more sense, well thx again

OpenStudy (anonymous):

Do you see how I arrived at \(2[u^2-2]\)?

OpenStudy (anonymous):

Cause I kinda skipped some steps there.

OpenStudy (anonymous):

well no, but I understand that it's the right equation to the integral

OpenStudy (anonymous):

I can show the steps.

OpenStudy (anonymous):

there are steps to it?

OpenStudy (anonymous):

Yes certainly.

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

Since \( u = (x+2)^{1/2} \implies x = u^2 -2\) \(\implies dx = 2u\ du\) \[\implies \frac{x}{(x+2)^{1/2}}dx = \frac{u^2-2}{u} *2u\ du\] \[=2[u^2-2]\ du\]

OpenStudy (anonymous):

awesome thanks a lot, I don't think we even learned that yet

OpenStudy (anonymous):

do you always implicitly differentiate x? (x'=dx and not 1?)

OpenStudy (anonymous):

I found that out nvm